OpenStudy (cggurumanjunath):

PROVE\quad THAT\quad 2{ n }^{ 2 }\ge { (n+1) }^{ 2 }\quad for\quad n\ge 3\quad mathematical\quad analysis\quad problem.

OpenStudy (cggurumanjunath):

\[PROVE\quad THAT\quad 2{ n }^{ 2 }\ge { (n+1) }^{ 2 }\quad for\quad n\ge 3\quad mathematical\quad analysis\quad problem.\]

terenzreignz (terenzreignz):

I'm tempted to say induction... are we allowed?

OpenStudy (cggurumanjunath):

this is a mathematical analysis problem.

OpenStudy (cggurumanjunath):

yes it is, induction is allowed.

terenzreignz (terenzreignz):

Well, first, show that it holds for n = 3.

OpenStudy (cggurumanjunath):

yes i have done that i got 18>16.

terenzreignz (terenzreignz):

Now, assume it holds for n = k, where k is some integer greater than or equal to 3. That is to say, this statement \[\Large 2k^2 \ge (k+1)^2 \qquad k\ge3\] is true.

OpenStudy (cggurumanjunath):

ok.

OpenStudy (cggurumanjunath):

now we have to show that it is true for (k+1) term right ?

terenzreignz (terenzreignz):

that's right :) Try it ^_^

OpenStudy (cggurumanjunath):

we get 2